How much Ione invest if she wants to have P19,640.62 at the end of 2.75 years at a bank that offers 11(1/4)% simple interest


1
Expert's answer
2021-12-14T16:02:38-0500

"A=P(1+rt)"


"19,640.62=P(1+\frac{45}{400}\times2.75)"


"19,640.62=1.309375P"


"P=\frac{19640.2}{1.309375}=P14,999.7\approx P15,000"